# Embracing Radiance Field Rendering in 6G: Over-the-Air Training and Inference with 3D Contents

### v1

@@ Abstract
The efficient representation, transmission, and reconstruction of three-dimensional (3D) contents are becoming increasingly important for sixth-generation (6G) networks that aim to merge virtual and physical worlds for offering immersive communication experiences. Neural radiance field (NeRF) and 3D Gaussian splatting (3D-GS) have recently emerged as two promising 3D representation techniques based on radiance field rendering, which are able to provide photorealistic rendering results for complex scenes. Therefore, embracing NeRF and 3D-GS in 6G networks is envisioned to be a prominent solution to support emerging 3D applications with enhanced quality of experience. This paper provides a comprehensive overview on the integration of NeRF and 3D-GS in 6G. First, we review the basics of the radiance field rendering techniques, and highlight their applications and implementation challenges over wireless networks. Next, we consider the over-the-air training of NeRF and 3D-GS models over wireless networks by presenting various learning techniques. We particularly focus on the federated learning design over a hierarchical device-edge-cloud architecture, which is suitable for exploiting distributed data and computing resources over 6G networks to train large models representing large-scale scenes. Then, we discuss three practical rendering architectures of NeRF and 3D-GS models at wireless network edge. We provide model compression approaches to facilitate the transmission of radiance field models, and present rendering acceleration approaches and joint computation and communication designs to enhance the rendering efficiency. In particular, we propose a new semantic communication enabled 3D content transmission design, in which the radiance field models are exploited as the semantic knowledge base to reduce the communication overhead for distributed inference. Furthermore, we present the utilization of radiance field rendering in wireless applications like radio mapping and radio imaging, in which radiance field models are used to effectively represent complex radio environments to facilitate wireless network designs. It is our hope that this paper can provide new insights on the interesting wireless integration with radiance field rendering for future 6G networks with 3D contents.
@@ I Introduction
Sixth-generation (6G) networks are experiencing a paradigm shift from the connected everything in fifth-generation (5G) to the new vision of connected intelligence, bridging the virtual and physical worlds. On the one hand, 6G networks need to utilize new wireless technologies such as millimeter (mmWave)/terahertz (THz) and extremely large-scale antenna arrays to provide ultra-high-data-rate, hyper-reliable, and low-latency communications. On the other hand, 6G networks are expected to evolve towards new multi-functional networks by integrating various functionalities such as wireless sensing, communication, mobile edge computation (MEC), and artificial intelligence (AI) [2, 3, 4]. The multi-functional operation thus helps meet the diverse requirements of emerging applications such as extended reality (XR), metaverse, autonomous driving, and intelligent robotics.
@@ II-A1 NeRF
In summary, the implicit representation of NeRF is able to represent 3D contents at a small data size and render realistic images without deteriorating details of 3D contents, but it requires the frequent querying of neural networks for volume rendering and thus is computationally inefficient. By contrast, the explicit representation of 3D-GS is intuitive and easy to edit and can render high-quality images without relying on neural networks or volume rendering, but it requires a large amount of data for representation, transmission, and storage. Therefore, there generally exists a tradeoff between the communication and computation in selecting the 3D representation approaches when implementing over wireless networks.
@@ II-B1 Wireless immersive video streaming
Radiance field rendering also holds significant potential to enhance 3D-based vision intelligence tasks such as video analytics [20], simultaneous localization and mapping (SLAM) [21], and robot navigation [22]. This is primarily attributed to the role of spatial information provided by 3D contents, which significantly enhances the inference capabilities required for these tasks. For instance, NeRF/3D-GS models can reconstruct high-quality scenes containing terrain information to facilitate video object detection and robot navigation. These tasks are implemented distributedly over 6G wireless networks, leveraging distributed computation and communication resources. For example, in a robot navigation task, the robot can transmit multi-view images to the edge server for training NeRF/3D-GS models, and subsequently, the edge server can utilize the model to render novel viewing images of regions that have not yet been traversed. These images can be transmitted back to the robot to assist global path planning and decision making. In these applications, the training and inference of radiance fields at the network edge need to be designed to properly exploit the distributed computation and communication resources.
@@ II-B3 Wireless sensing and environment-aware communications
Motivated by the great success in the field of optical imagery, radiance field rendering is expected to be applicable in the wireless field, especially for the representation of wirelessly generated 3D information by effectively modeling the physical process of wireless signal propagation. This application is becoming more and more important for 6G networks, as radio sensing and imaging are seamlessly integrated as a new function of wireless networks with the advancement of integrated sensing and communication (ISAC) [4, 23]. In particular, radiance field rendering such as NeRF can be used to model the radio frequency (RF) signal attenuation process including signal reflection, scattering, and refraction, and accordingly represent radio maps that store the channel information at different transceiver locations [24]. In addition, radiance field rendering is also useful for radar imaging (e.g., via synthetic aperture radar (SAR)) [25], in which the radio imaging process can be modeled by using NeRF, such that the SAR imaging information can be obtained from different views. Furthermore, radiance field rendering can benefit sensing-assisted environment-aware communications [26], in which radiance field models can be used to represent the 3D environment with considerable visual quality and multi-view vision information, such that the environmental obstacles and scatterers can be detected to infer and predict the channel knowledge information [27, 28].
@@ III-B Hierarchical Federated Radiance Field Learning
Due to the benefits in exploiting distributed computation power and preserving data privacy, federated radiance field learning is particularly appealing for training over wireless networks. Nevertheless, prior works on federated radiance field learning focused on designing the learning algorithm, without considering specific wireless network architectures. In this paper, we present a new device-edge-cloud hierarchical federated learning architecture for the training of radiance field models for a large-scale scene. Specifically, this architecture consists of one cloud server, $L>1$ edge servers, and a large number of $K>L$ end devices, as shown in Fig. 3. At the lower level of the hierarchical architecture, the end devices are each installed with one or more cameras to capture the multi-view observation data, and the edge servers each cover a certain area to coordinate multiple associated end devices. At the higher level, the cloud center coordinates the training operation of these edge servers. The device-edge-cloud hierarchical federated radiance field learning approach is appealing for the efficient training of large-scale scene radiance fields (e.g., at a city level). In particular, the hierarchical architecture allows the cloud center to coordinate massive end devices through a limited number of edge servers, thus significantly reducing the traffic loads over the whole network. Next, each end devices only need to send their local model information to nearby edge servers for aggregating edge models of a relatively smaller-scale sub-scene, thus enhancing the training efficiency. Furthermore, the cloud center can aggregate the edge models from edge servers to update the global models for the large-scale scene. In addition, the cloud center can also adjust the learning parameters (e.g., frequencies of edge and global aggregations) and coordinate the network-level resource management, by fully exploiting the diverse communication, computation, and storage powers over the whole network.
@@
In particular, the hierarchical federated radiance field learning is implemented as follows. Let $D_{k}=\{(I_{k,i},p_{k,i})\}$ denote the local dataset at end device $k,~{}k\in\{1,\ldots,K\}$, where $I_{k,i}\in\mathbb{R}^{H\times W\times 3}$ and $p_{k,i}\in\mathbb{R}^{2}$ denote the $i$-th observed images and the corresponding viewing direction by device $k$, respectively. Here, $H$ and $W$ denote the height and width of the observed image, respectively. Due to the inherent limitations in camera views, each device can only access the multi-view observation data from a subsection of the entire scene. Under this setup, the training of radiance field models is implemented over nested two rounds of iterations. In each lower-level or inner iteration, the end devices each perform local training (via multiple rounds of stochastic gradient descent (SGD) calculations) based on the captured data and then upload their local models and the 3D location information to their associated edge servers. After receiving the local models and local location information, each edge server aligns the pose of its associated devices in a common coordinate, and then aggregates the received local models to obtain updated edge models. In each upper-level or outer iteration, the edge servers transmit the updated edge models (after one or more rounds of inner iterations) and the corresponding coordinate information to the cloud center, such that the cloud center can align the coordinates and then aggregate the edge models to obtain the global models. In general, the frequencies of inner and outer iterations (i.e., the number of local SGD calculations within each inner iterations and the number of inner iterations or edge aggregations within each outer iteration) may affect the consumed computation and communication resources towards convergence, and thus are key decision variables for performance optimization.
@@
Notice that the pose alignment presents a unique challenge in federated radiance field learning for large-scale scenarios, where end devices (such as UAVs) are dispersed across various locations, each observing the scene from unique viewing directions [33]. In this case, the local models obtained at different end devices have distinct poses, which thus cannot be directly aggregated at the edge server. Instead, the edge server needs to first use proper pose alignment to revise these poses towards the common coordinate, and then perform the edge aggregation. As such, it is critical for the end devices (and edge servers) to upload their pose information (and coordinate information) to facilitate the edge (and cloud) aggregation. Some useful device pose alignment methods can be referred to in, e.g., [33, 34].
@@ III-D3 Vertical Federated Radiance Field Learning
Vertical federated radiance field learning is another interesting issue, where datasets from different devices share the same sample space but differ in feature spaces [45]. This is the case when depth cameras are practically involved to generate depth maps (reflecting spatial information of scenes), in addition to the conventional cameras generating multi-view RGB images. To leverage the overall data collected from RGB and depth cameras for efficient radiance field model learning, we should consider vertical federated learning paradigm together with depth images based radiance field construction algorithms (see, e.g., [46]).
@@ IV-A3 Co-Rendering
To fully utilize the computation resources at edges and devices as well as exploit both benefits of local and edge rendering, device-edge co-inference or co-rendering has emerged as a promising solution, which has been widely adopted in various AI applications nowadays [48]. In the device-edge co-inference or local-remote co-rendering, the radiance field models (e.g., the MLP neural network for NeRF) are split into two parts to be deployed at end devices and edge servers, respectively, such that the rendering is implemented distributedly at the two types of nodes. In the co-rendering process, the end devices first obtain the opacity and color of sampled points along the rays based on the deployed partial radiance field models (e.g., MLPs), which are then sent to the server for volume rendering. Finally, the server sends the rendering results back to the devices. The co-rendering design is expected to enhance the rendering performance via properly designing the task allocation between the devices and servers, for which the joint communication and computation designs become essential.
@@ IV-B1 Model Compression
Compressing the radiance field models is essential to their efficient transmission for model deployment, especially when the training nodes generating the models are far-apart from the end users and the size of trained models is big for large-scale scenes. Efficient radiance field model compression approaches generally include network pruning, quantization, low-rank approximation, and knowledge distillation [16, 17]. For NeRF, network pruning and weight quantization are efficient to remove the redundant network layers and transform the full-precision floating-point numbers into lower-bit representations, respectively. Furthermore, low-rank approximation is applicable to replace the high-rank matrices with low-rank ones, while knowledge distillation extracts knowledge from a large model and condenses it into a smaller one. Different from NeRF, the size of 3D-GS models depends on the volume of explicit data, i.e., the number of Gaussian functions. To efficiently compress 3D-GS models, pruning [49] and quantization [50] are promising to reduce the number of Gaussian functions for decreasing the amount of explicit data, while maintaining the rendering quality.

### v2

@@ Abstract
The efficient representation, transmission, and reconstruction of three-dimensional (3D) contents are becoming increasingly important for sixth-generation (6G) networks that aim to merge virtual and physical worlds for offering immersive communication experiences. Neural radiance field (NeRF) and 3D Gaussian splatting (3D-GS) have recently emerged as two promising 3D representation techniques based on radiance field rendering, which are able to provide photorealistic rendering results for complex scenes. Therefore, embracing NeRF and 3D-GS in 6G networks is envisioned to be a prominent solution to support emerging 3D applications with enhanced quality of experience. This paper provides a comprehensive overview on the integration of NeRF and 3D-GS in 6G. First, we review the basics of the radiance field rendering techniques, and highlight their applications and implementation challenges over wireless networks. Next, we consider the over-the-air training of NeRF and 3D-GS models over wireless networks by presenting various learning techniques. We particularly focus on the federated learning design over a hierarchical device-edge-cloud architecture, which is suitable for exploiting distributed data and computing resources over 6G networks to train large models representing large-scale scenes. Then, we consider the over-the-air rendering of NeRF and 3D-GS models at wireless network edge. We present three practical rendering architectures, namely local, remote, and co-rendering, respectively, and provide model compression approaches to facilitate the transmission of radiance field models for rendering. We also present rendering acceleration approaches and joint computation and communication designs to enhance the rendering efficiency. In a case study, we propose a new semantic communication enabled 3D content transmission design, in which the radiance field models are exploited as the semantic knowledge base to reduce the communication overhead for distributed inference. In addition, we discuss the utilization of radiance field rendering in wireless applications like radio mapping and radio imaging, in which radiance field models are used to effectively represent complex radio environments to facilitate wireless network designs. It is our hope that this paper can provide new insights on the interesting wireless integration with radiance field rendering for future 6G networks with 3D contents.
@@ I Introduction
Sixth-generation (6G) networks are experiencing a paradigm shift from the connected everything in fifth-generation (5G) to the new vision of connected intelligence, bridging the virtual and physical worlds. On the one hand, 6G networks are envisioned to utilize new wireless technologies such as millimeter wave (mmWave)/terahertz (THz) and extremely large-scale antenna arrays, which can provide ultra-high-data-rate, hyper-reliable, and ultra-low-latency communications to handle the massive data transmission demands. On the other hand, 6G networks are expected to evolve towards new multi-functional networks by integrating various functionalities such as wireless sensing, communication, mobile edge computation (MEC), and artificial intelligence (AI) [2, 3, 4]. The multi-functional operation facilitates the fusion of emerging AI technologies and wireless networks, and thus helps meet the diverse requirements of emerging applications such as extended reality (XR), metaverse, autonomous driving, and intelligent robotics.
@@ II-A1 NeRF
In summary, the implicit representation of NeRF is able to represent 3D contents by a lightweight model and render realistic images without deteriorating details of 3D contents, but it requires the frequent querying of neural networks for volume rendering and thus is computationally inefficient. By contrast, the explicit representation of 3D-GS is intuitive and easy to edit and can render high-quality images without relying on neural networks or volume rendering and enables parallel rendering processes, but it requires a large amount of data for representation, transmission, and storage. Therefore, there generally exists a tradeoff between the communication, computation, and storage in selecting the 3D representation approaches when implementing radiance field rendering in diver scenarios over wireless networks.
@@ II-B1 Wireless immersive video streaming
Radiance field rendering also holds significant potential to enhance 3D-based vision intelligence tasks such as video analytics [20], simultaneous localization and mapping (SLAM) [21], and robot navigation [22]. This is primarily attributed to the role of spatial information provided by 3D contents, which significantly enhances the inference capabilities required for these tasks. For instance, NeRF/3D-GS models can reconstruct high-quality scenes containing terrain information to facilitate video object detection and robot navigation. In the 6G era, these tasks are implemented at wirelessly connected agent/robots and remote servers in a distributed manner [23, 24], by fully leveraging their distributed computation and communication resources. For example, in a robot navigation task, the robots can transmit their captured multi-view images to the edge server for training NeRF/3D-GS models, based on which, the edge server can subsequently render images of novel viewing regions by leveraging the rich computation resource. Next, the server can transmit these novel-viewing images back to the robots to assist their global path planning and decision making. In these applications, the training and inference of radiance fields at the network edge need to be designed by properly exploiting the distributed computation and communication resources.
@@ II-B3 Wireless sensing and environment-aware communications
Motivated by the great success in the field of optical imagery, radiance field rendering is expected to be applicable in the wireless field, especially for the representation of wirelessly generated 3D information by effectively modeling the physical process of wireless signal propagation. This application is becoming increasingly important for 6G networks, as radio sensing and imaging are seamlessly integrated as a new function of wireless networks with the advancement of integrated sensing and communication (ISAC) [4, 25, 26]. Radiance field rendering can benefit the fusion of sensing and communication functionalities by enhancing the sensing performance with 3D spatial information and provides multi-modal information to facilitate the communication designs. In particular, radiance field rendering such as NeRF can be used to model the radio frequency (RF) signal attenuation process including signal reflection, scattering, and refraction, and accordingly represent radio maps that store the channel information at different transceiver locations [27]. In addition, radiance field rendering is also useful for radar imaging (e.g., via synthetic aperture radar (SAR)) [28], in which the radio imaging process can be modeled by using NeRF, such that the SAR imaging information can be obtained from different views. Furthermore, radiance field rendering can benefit sensing-assisted environment-aware communications [29], in which radiance field models can be used to represent the 3D environment with considerable visual quality and multi-view vision information, such that the environmental obstacles and scatterers can be detected to infer and predict the channel knowledge information [30, 31].
@@ III-A2 Distributed (Federated) Radiance Field Learning
In summary, the centralized radiance field learning is able to process the multi-view data and train the radiance field model in a central network node without coordinating distributed devices, but it may incur large communication overheads and privacy leakage issues. By contrast, the federated radiance field learning aims at exploiting the distributed computation resources among separate devices to train the models without sharing raw data, which, however, generally requires large iterations of model exchange between devices and servers.
@@ III-B Hierarchical Federated Radiance Field Learning
Due to the benefits in exploiting distributed computation power and preserving data privacy, federated radiance field learning is particularly appealing for training over wireless networks. Nevertheless, prior works on federated radiance field learning focused on designing the learning algorithm, without considering specific wireless network architectures. In this paper, we present a new device-edge-cloud hierarchical federated learning architecture for the training of radiance field models for a large-scale scene.111 Generating 3D contents for large-scale scenes may have abundant applications in various 6G applications like low-altitude economy [36], smart city [37], and intelligent transportation [24]. The design principles for large-scale scenes are also applicable to the case with small-scale scenes [38]. Specifically, this architecture consists of one cloud server, $L>1$ edge servers, and a large number of $K>L$ end devices, as shown in Fig. 3. At the lower level of the hierarchical architecture, the end devices are each installed with one or more cameras to capture the multi-view observation data, and the edge servers each cover a certain area to coordinate multiple associated end devices. At the higher level, the cloud center coordinates the training operation of these edge servers. The device-edge-cloud hierarchical federated radiance field learning approach is appealing for the efficient training of large-scale scene radiance fields (e.g., at a city level). In particular, the hierarchical architecture allows the cloud center to coordinate massive end devices through a limited number of edge servers, thus significantly reducing the traffic loads over the whole network. Next, each end devices only need to send their local model information to nearby edge servers for aggregating edge models of a relatively smaller-scale sub-scene, thus enhancing the training efficiency. Furthermore, the cloud center can aggregate the edge models from edge servers to update the global models for the large-scale scene. In addition, the cloud center can also adjust the learning parameters (e.g., frequencies of edge and global aggregations) and coordinate the network-level resource management, by fully exploiting the diverse communication, computation, and storage powers over the whole network.
To begin with, we briefly introduce the dataset for training the radiance field models. Let $D_{k}=\{(I_{k,i},p_{k,i})\}$ denote the local dataset at device $k\in\{1,\ldots,K\}$, where $I_{k,i}\in\mathbb{R}^{H\times W\times 3}$ denotes the $i$-th observed image by device $k$ that serves as the ground truth during the training process. Here, $H$ and $W$ denote the height and width of the observed image, respectively. Furthermore, $p_{k,i}$ denotes the corresponding viewing information, which contains the extrinsic matrix and intrinsic matrix of the cameras. The extrinsic matrix exploits the rotation and translation information to transform world coordinates into camera coordinates, while its inverse matrix can convert camera coordinates back to world coordinates. The intrinsic matrix uses the focal information to project camera coordinates onto the 2D image plane. In computer vision society, there are two methods to obtain the dataset. The first method is to directly extract images and the corresponding viewing information from synthetic datasets [12]. The second method is to utilize the COLMAP structure-from-motion package [39] to estimate the extrinsic and intrinsic matrices for real observed images [40].
@@
Then, the hierarchical federated radiance field learning is implemented as follows by particularly considering the use of 3D-GS models. Due to the inherent limitations in camera views, each device can only access the multi-view observation data from a subsection of the entire scene. Under this setup, the training of 3D-GS models is implemented over nested two rounds of iterations. Specifically, we denote the local 3D-GS model at device $k$ as $G_{\text{local}}^{k}=\{\mathcal{G}^{(i)}\}=\{\hat{\boldsymbol{x}}^{(i)},\Sigma^{(i)},\alpha^{(i)},c^{(i)}\}_{i}$, which is a set of 3D Gaussians. Similarly, we denote the edge 3D-GS model at edge server $l$ as $G_{\text{edge}}^{l}=\{\hat{\boldsymbol{x}}^{(i)},\Sigma^{(i)},\alpha^{(i)},c^{(i)}\}_{i}$ and the global model at cloud center as $G_{\text{global}}$. In each lower-level or inner iteration, the end devices each perform local training (via multiple rounds of stochastic gradient descent (SGD) calculations) based on the captured data and then upload their local models and the 3D location information to their associated edge servers. After receiving the local models and local location information, each edge server aligns the pose of its associated devices in a common coordinate, and then aggregates the received local models to obtain updated edge models. In particular, each edge server $l$ obtain the edge model via merging the local models as $G_{\text{edge}}^{l}=G_{\text{local}}^{1}\cup G_{\text{local}}^{2}\cup,\ldots,\cup~{}G_{\text{local}}^{\hat{k}}$, where $\hat{k}$ denotes the number of associated end devices of edge server $l$. Then, in each upper-level or outer iteration, the edge servers transmit the updated edge models (after one or more rounds of inner iterations) and the corresponding coordinate information to the cloud center, such that the cloud center can align the coordinates and then aggregate the edge models to obtain the global models. Similarly, the cloud center obtains the global model via merging the edge models as $G_{\text{global}}=G_{\text{edge}}^{1}\cup G_{\text{edge}}^{2}\cup,\ldots,\cup~{}G_{\text{edge}}^{L}$. In general, the frequencies of inner and outer iterations (i.e., the number of local SGD calculations within each inner iterations and the number of inner iterations or edge aggregations within each outer iteration) may affect the consumed computation and communication resources towards convergence, and thus are key decision variables for performance optimization.
@@
Notice that the pose alignment presents a unique challenge in federated radiance field learning for large-scale scenarios, where end devices (such as UAVs) are dispersed across various locations, each observing the scene from unique viewing directions and thus obtaining the multi-view dataset with different pose information $p_{k,i}$ [41, 42]. In this case, the local models obtained at different end devices have distinct poses, which cannot be directly aggregated at the edge server. As such, the edge server needs to first use proper pose alignment to revise these poses towards the common coordinate, and then perform the edge aggregation. As such, it is critical for the end devices (and edge servers) to upload their pose information (and coordinate information) to facilitate the edge (and cloud) aggregation. Some useful device pose alignment methods can be referred to in, e.g., [41, 42].
@@ III-D3 Vertical Federated Radiance Field Learning
Vertical federated radiance field learning is another interesting issue, where datasets from different devices share the same sample space but differ in feature spaces [53]. This is the case when depth cameras are practically involved in generating depth maps (reflecting spatial information of scenes), in addition to the conventional cameras generating multi-view RGB images. To leverage the overall data collected from RGB and depth cameras for efficient radiance field model learning, we should consider vertical federated learning paradigm together with depth images based radiance field construction algorithms (see, e.g., [54]).
@@ III-D4 Over-the-air Radiance Field Federated Learning
III-D5 Radiance Field Federated Learning for Dynamic Scenes
Dynamic scene representation is an important issue in radiance field rendering, which has been investigated in prior works [56, 57, 58]. In general, there are two methods to represent the dynamic scenes. Firstly, for NeRF, deformable neural networks are widely adopted to represent dynamic scenes. For instance, the work [56] embeds an additional time dimension into the model to control the color and density for time-varying scenes. This allows the model to account for dynamic changes over time. For 3D-GS, deformable fields are exploited to dynamically control the position and 3D covariance of Gaussians over time [57]. This approach enables the 3D-GS model to adjust dynamically as the scene changes, providing accurate representations of dynamic scenes. Inspired by these methods, we can employ federated learning to train deformable radiance field models to represent dynamic scenes. Furthermore, another promising method for dynamic scene representation is to first train a generalized radiance field model to predict coarse color and density information and then fine-tune it to predict finer details for accurately rendering dynamic scenes [58].
@@ IV-A3 Co-Rendering
To fully utilize the computation resources at edges and devices as well as exploit both benefits of local and edge rendering, device-edge co-inference or co-rendering has emerged as a promising solution, which has been widely adopted in various AI applications nowadays [59]. In the device-edge co-inference or local-remote co-rendering, the radiance field models (e.g., the MLP neural network for NeRF) are split into two parts to be deployed at end devices and edge servers, respectively, such that the rendering is implemented distributedly at the two types of nodes. In the co-rendering process, the end devices first obtain the opacities and colors of sampled points along the rays based on the deployed partial radiance field models (e.g., MLPs). These colors and opacities serve as the intermediate parameters, which are then sent to the server for volume rendering. Finally, the server sends the rendering results back to the devices. The co-rendering design is expected to enhance the rendering performance via properly designing the task allocation between the devices and servers, for which the joint communication and computation designs become essential.
In summary, the local rendering, edge rendering, and co-rendering have their pros and cons in terms of rendering latency and computation/storage resource requirements. First, local rendering deploys trained radiance field models on local devices, enabling instant rendering by leveraging local computational resources and avoiding long-distance transmission. This approach minimizes latency but demands lightweight models due to limited local computational and storage capacities. Next, edge rendering allows the end devices to send viewing information to edge servers, which render images and then send the rendering results back. This may result in substantial latency due to two-way communication, but can handle larger models due to the richer computational resources at edge servers. Furthermore, co-rendering combines the strengths of local and edge rendering by utilizing distributed resources of end devices while offloading heavy tasks to edge servers, thus improving the overall rendering performance. Despite its efficiency, co-rendering may still encounter latency issues from the intermediate parameter exchange between devices and servers. Therefore, the choice of proper rendering architecture should consider specific latency requirements and available computational/storage resources.
@@ IV-B1 Model Compression
Compressing the radiance field models is essential for model deployment, especially when the training nodes generating the models are far-apart from the end users and the size of trained models becomes large, e.g., for representing large-scale scenes. The model compression can reduce the model size, ensuring faster and more reliable transmission, which is particularly important in scenarios with limited bandwidth or under stringent latency requirements. Efficient radiance field model compression approaches generally include network pruning, quantization, low-rank approximation, and knowledge distillation [16, 17]. For NeRF, network pruning and weight quantization are efficient to remove the redundant network layers and transform the full-precision floating-point numbers into lower-bit representations, respectively. Furthermore, low-rank approximation is applicable to replace the high-rank matrices with low-rank ones, while knowledge distillation extracts knowledge from a large model and condenses it into a smaller one. Different from NeRF, the size of 3D-GS models depends on the volume of explicit data, i.e., the number of Gaussian functions. To efficiently compress 3D-GS models, pruning [60] and quantization [61] are promising to reduce the number of Gaussian functions for decreasing the amount of explicit data, while maintaining the rendering quality.